# DitHub: A Modular Framework for Incremental Open-Vocabulary Object Detection

### v1

@@ B.1 ODinW-13 and MS COCO
Unlike structured and standardized datasets such as MS COCO, ODinW-13 aggregates data from thirteen sources: Aerial Maritime Drone (Ae), Aquarium (Aq), Cottontail Rabbit (Co), EgoHands (Eg), Mushrooms (Mu), Packages (Pa), Pascal VOC (Pv), Pistols (Pi), Pothole (Po), Raccoon (Ra), Shellfish (Sh), Thermal Dogs and People (Th), Vehicles (Ve). This comprehensive collection covers rare object categories, diverse contextual conditions, and challenging domains. Consequently, the ODinW-13 benchmark offers an ideal scenario to evaluate the capability of DitHub to sequentially adapt to heterogeneous task distributions while effectively mitigating catastrophic forgetting on previously learned tasks.
@@
B.2 ODinW-O
We introduce a new benchmark, ODinW-O (Overlapped), designed to further evaluate the class-incremental and specialization capabilities of DitHub. Derived from ODinW-35, ODinW-O retains only classes shared across at least one task (overlapping classes), while removing tasks with no shared classes. This filtering is necessary because classes that do not reappear in subsequent tasks present an ideal case in the DitHub framework. Specifically, their corresponding non-recurring class modules remain frozen, requiring no further adaptation. In contrast, ODinW-O eliminates this scenario by ensuring that all retained classes appear in multiple tasks. By enforcing strictly overlapping classes, ODinW-O provides a challenging benchmark for evaluating incremental learning, specifically tailored to assess the robustness of our method. Additionally, to rigorously evaluate the model robustness to domain shifts, we designed ODinW-O by explicitly selecting datasets that feature markedly different contexts (Fig. B), enabling clear insights into how effectively the model adapts to variations in visual appearance, environmental conditions, and imaging domains. We select the following datasets, already presented in Sec. B.1, and extracted only the overlapping classes:
  • Aerial Maritime Drone: boat and car,
  • Pascal VOC: boat, car, dog, and person
  • Thermal Dogs and People: dog and person,
  • Vehicles: car and truck,
@@ B.2 ODinW-O
  • Hard Hat Workers: person
  • Self Driving Cars: car and truck
@@
Person Car
Refer to caption Refer to caption
Refer to caption Refer to caption
Appendix C Implementation Details
All experiments for DitHub were conducted on a single RTX A5000 GPU with 24GB of VRAM, requiring approximately 8 hours per training run with a batch size of 2. During training, our implementation in PyTorch efficiently utilizes around 10 GB of VRAM. We employed GroundingDINO-T (Swin-T) as an open-vocabulary object detector, pretrained on O365 [39], GoldG [17], and Cap4M [17] datasets. We performed training with random shuffling across three different seeds to ensure robustness.
For LoRA fine-tuning, we perform low-rank updates to the attention layers with a rank of 16, applying a learning rate of $1\times 10^{-3}$ and a weight decay of $1\times 10^{-2}$. This approach enables efficient adaptation while significantly reducing computational overhead and memory storage for each class-specialized module. Optimization is performed using AdamW. We rely solely on the contrastive classification loss and the localization loss of Grounding DINO, without introducing any additional loss functions. The full Grounding DINO model, after integrating 44 classes during training on ODinW-13, consists of approximately 173 million parameters, with 19.6 million LoRA parameters actively updated during training. After fine-tuning on ODinW-13, only 75.65 MB of LoRA parameters are retained for storage.
@@ Appendix D Additional Results
Ours ±$\pm 0.24$$\pm 0.66$$\pm 2.39$$\pm 1.88$$\pm 0.76$$\pm 1.31$$\pm 0.46$$\pm 0.84$
Table C: Standard deviations of the reproduced methods in Table 2.
D.1 Standard Deviations
Shots Method ZCOCO Avg Ae Aq Co Eg Mu Pa Pv Pi Po Ra Sh Th Ve
1 ZiRa $\pm$ 0.90 $\pm$ 2.05 $\pm$ 1.49 $\pm$ 0.17 $\pm$ 2.16 $\pm$ 0.98 $\pm$ 0.79 $\pm$ 1.63 $\pm$ 1.33 $\pm$ 7.91 $\pm$ 3.29 $\pm$ 12.17 $\pm$ 4.45 $\pm$ 3.07 $\pm$ 3.25
DitHub $\pm$ 0.60 $\pm$ 0.81 $\pm$ 0.95 $\pm$ 0.57 $\pm$ 2.66 $\pm$ 1.58 $\pm$ 3.72 $\pm$ 0.00 $\pm$ 1.53 $\pm$ 1.10 $\pm$ 2.90 $\pm$ 7.03 $\pm$ 3.68 $\pm$ 4.66 $\pm$ 2.46
5 ZiRa $\pm$ 0.23 $\pm$ 0.43 $\pm$ 3.76 $\pm$ 2.15 $\pm$ 1.40 $\pm$ 1.08 $\pm$ 3.54 $\pm$ 3.74 $\pm$ 1.54 $\pm$ 4.26 $\pm$ 3.73 $\pm$ 4.13 $\pm$ 2.03 $\pm$ 2.02 $\pm$ 4.16
DitHub $\pm$ 0.42 $\pm$ 1.27 $\pm$ 2.52 $\pm$ 0.70 $\pm$ 2.42 $\pm$ 6.74 $\pm$ 0.76 $\pm$ 3.27 $\pm$ 0.17 $\pm$ 5.49 $\pm$ 4.79 $\pm$ 2.22 $\pm$ 4.29 $\pm$ 1.81 $\pm$ 1.54
10 ZiRa $\pm$ 0.28 $\pm$ 0.66 $\pm$ 0.44 $\pm$ 2.64 $\pm$ 2.77 $\pm$ 0.52 $\pm$ 6.43 $\pm$ 0.97 $\pm$ 1.04 $\pm$ 2.17 $\pm$ 2.04 $\pm$ 3.90 $\pm$ 1.21 $\pm$ 3.71 $\pm$ 0.87
DitHub $\pm$ 0.20 $\pm$ 0.54 $\pm$ 2.26 $\pm$ 0.72 $\pm$ 4.23 $\pm$ 3.53 $\pm$ 1.35 $\pm$ 3.19 $\pm$ 0.30 $\pm$ 2.74 $\pm$ 0.34 $\pm$ 2.99 $\pm$ 1.01 $\pm$ 1.15 $\pm$ 1.27
Full ZiRA $\pm$ 0.44 $\pm$ 0.68 $\pm$ 2.73 $\pm$ 1.17 $\pm$ 1.39 $\pm$ 4.77 $\pm$ 7.26 $\pm$ 2.19 $\pm$ 1.33 $\pm$ 0.33 $\pm$ 4.14 $\pm$ 3.57 $\pm$ 2.44 $\pm$ 2.82 $\pm$ 3.13
DitHub $\pm$ 0.16 $\pm$ 0.44 $\pm$ 2.39 $\pm$ 1.00 $\pm$ 2.83 $\pm$ 0.48 $\pm$ 1.34 $\pm$ 0.60 $\pm$ 0.38 $\pm$ 1.68 $\pm$ 0.76 $\pm$ 1.57 $\pm$ 2.87 $\pm$ 0.69 $\pm$ 1.01
Table D: Standard deviations of the reproduced methods in Table 1.
We report the standard deviations of the reproduced methods from Tab. 1 and Tab. 2 in Tab. D and C, respectively, computed over three runs with different random seeds.
D.2 Unlearning Additionals
Figure C: Matrix visualization of mAP degradation per class after removing class modules. Colors are normalized column-wise for readability; darker cells indicate greater performance drops.
As discussed in Sec. 5.2, our modular approach enables the precise removal of specific knowledge from Grounding DINO by subtracting corresponding experts. Figure C illustrates this process: each column represents a target class’s mAP after unlearning different modules, with colors normalized column-wise to highlight relative drops. The darker diagonal pattern indicates that removing a class module leads to the most significant performance drop for its corresponding class. This confirms that our modules effectively encapsulate specialized knowledge, as unlearning them selectively impacts their respective classes.
Appendix E Qualitatives
@@
In Fig. D, we compare zero-shot Grounding DINO (left) with ZiRA (center) and DitHub (right). Our method excels at detecting challenging objects, including small ones that other models miss. Additionally, DitHub avoids a false positive (ZiRA) and a false negative (Grounding DINO), demonstrating more reliable predictions. It also correctly identifies both cars (last row) despite occlusions and distractors. This improved stability stems from per-class specialization, which mitigates hallucinations and ensures smoother adaptation to domain shifts.

### v2

@@ C.1 ODinW-13 and MS COCO
Unlike structured and standardized datasets such as MS COCO, ODinW-13 aggregates data from thirteen sources: Aerial Maritime Drone (Ae), Aquarium (Aq), Cottontail Rabbit (Co), EgoHands (Eg), Mushrooms (Mu), Packages (Pa), Pascal VOC (Pv), Pistols (Pi), Pothole (Po), Raccoon (Ra), Shellfish (Sh), Thermal Dogs and People (Th), Vehicles (Ve). This comprehensive collection covers rare object categories, diverse contextual conditions, and challenging domains. Consequently, the ODinW-13 benchmark offers an ideal scenario to evaluate the capability of DitHub to sequentially adapt to heterogeneous task distributions.
@@
C.2 ODinW-O
Person Car Truck Dog Boat
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
We introduce a new benchmark, ODinW-O (Overlapped), designed to further evaluate the class-incremental and specialization capabilities of DitHub. Derived from ODinW-35 [19], ODinW-O retains only classes shared across at least one task (i.e., overlapping classes), while removing tasks with no shared classes. This filtering is crucial because retaining classes that do not reappear in subsequent tasks would cause the benchmark to resemble the original one. Specifically, the corresponding non-recurring class modules would remain frozen, requiring no further adaptation. In contrast, ODinW-O eliminates this scenario by ensuring all retained classes appear in multiple tasks. By enforcing strictly overlapping classes, ODinW-O provides a challenging benchmark for evaluating Incremental Learning. Additionally, to rigorously evaluate the model robustness to domain shifts, we designed ODinW-O by explicitly selecting datasets that feature markedly different contexts (Fig. B), enabling clear insights into how effectively the model adapts to variations in visual appearance, environmental conditions, and imaging domains. We select the following datasets, already mentioned in Sec. C.1, and extracted only the overlapping classes:
  • Aerial Maritime Drone: boat and car;
  • Pascal VOC: boat, car, dog, and person;
  • Thermal Dogs and People: dog and person;
  • Vehicles: car and truck.
@@ C.2 ODinW-O
  • Hard Hat Workers: person;
  • Self Driving Cars: car and truck.
@@ Appendix D Additional Results
Figure C: Per-class mAP degradation after removing class modules. Darker cells indicate greater performance drops.
D.1 Standard Deviations
We report the standard deviations of the reproduced methods from Tab. 1 and Tab. 2 in Tab. E and F, respectively, computed over three runs with different random seeds.
D.2 Unlearning Additionals
As previously discussed in Sec. 5.2, our modular framework facilitates the precise removal of specific knowledge from Grounding DINO through the subtraction of corresponding experts. Fig. C visually represents this process: each column depicts the mean Average Precision (mAP) of a target class after unlearning different modules (reported on the y-axis), with colors normalized column-wise to emphasize relative performance decrements. The prominent darker diagonal pattern signifies that removing a class-specific module results in the most substantial performance reduction for its associated class. This observation substantiates that our modules effectively encapsulate specialized knowledge, as their unlearning selectively impacts their respective classes. Furthermore, this matrix representation elucidates correlations between certain classes during the unlearning procedure. For instance, the darker shading in the “bicycle” column indicates that the "person" class exhibits the second most significant mAP drop. This correlation arises from the frequent co-occurrence of “person” instances riding bicycles, naturally linking these classes within the comprehensive unlearning pipeline.
Table E: Standard deviations of the reproduced methods in Tab. 1.
Shots Method ZCOCO Avg Ae Aq Co Eg Mu Pa Pv Pi Po Ra Sh Th Ve
1 ZiRa ±$\pm 0.90$$\pm 2.05$$\pm 1.49$$\pm 0.17$$\pm 2.16$$\pm 0.98$$\pm 0.79$$\pm 1.63$$\pm 1.33$$\pm 7.91$$\pm 3.29$$\pm 12.17$$\pm 4.45$$\pm 3.07$$\pm 3.25$
DitHub ±$\pm 0.60$$\pm 0.81$$\pm 0.95$$\pm 0.57$$\pm 2.66$$\pm 1.58$$\pm 3.72$$\pm 0.00$$\pm 1.53$$\pm 1.10$$\pm 2.90$$\pm 7.03$$\pm 3.68$$\pm 4.66$$\pm 2.46$
5 ZiRa ±$\pm 0.23$$\pm 0.43$$\pm 3.76$$\pm 2.15$$\pm 1.40$$\pm 1.08$$\pm 3.54$$\pm 3.74$$\pm 1.54$$\pm 4.26$$\pm 3.73$$\pm 4.13$$\pm 2.03$$\pm 2.02$$\pm 4.16$
DitHub ±$\pm 0.42$$\pm 1.27$$\pm 2.52$$\pm 0.70$$\pm 2.42$$\pm 6.74$$\pm 0.76$$\pm 3.27$$\pm 0.17$$\pm 5.49$$\pm 4.79$$\pm 2.22$$\pm 4.29$$\pm 1.81$$\pm 1.54$
10 ZiRa ±$\pm 0.28$$\pm 0.66$$\pm 0.44$$\pm 2.64$$\pm 2.77$$\pm 0.52$$\pm 6.43$$\pm 0.97$$\pm 1.04$$\pm 2.17$$\pm 2.04$$\pm 3.90$$\pm 1.21$$\pm 3.71$$\pm 0.87$
DitHub ±$\pm 0.20$$\pm 0.54$$\pm 2.26$$\pm 0.72$$\pm 4.23$$\pm 3.53$$\pm 1.35$$\pm 3.19$$\pm 0.30$$\pm 2.74$$\pm 0.34$$\pm 2.99$$\pm 1.01$$\pm 1.15$$\pm 1.27$
Full ZiRA ±$\pm 0.44$$\pm 0.68$$\pm 2.73$$\pm 1.17$$\pm 1.39$$\pm 4.77$$\pm 7.26$$\pm 2.19$$\pm 1.33$$\pm 0.33$$\pm 4.14$$\pm 3.57$$\pm 2.44$$\pm 2.82$$\pm 3.13$
DitHub ±$\pm 0.16$$\pm 0.44$$\pm 2.39$$\pm 1.00$$\pm 2.83$$\pm 0.48$$\pm 1.34$$\pm 0.60$$\pm 0.38$$\pm 1.68$$\pm 0.76$$\pm 1.57$$\pm 2.87$$\pm 0.69$$\pm 1.01$
Table F: Standard deviations of the reproduced methods in Tab. 2.
@@ D.2 Unlearning Additionals
DitHub ±$\pm 0.24$$\pm 0.66$$\pm 2.39$$\pm 1.88$$\pm 0.76$$\pm 1.31$$\pm 0.46$$\pm 0.84$
Appendix E Implementation Details
All experiments for DitHub were conducted on a single RTX A5000 GPU with 24 GB of VRAM. Each training run, with a batch size of $2$, required approximately $8$ hours. Our PyTorch implementation efficiently utilized around 10 GB of VRAM during training. We employed Grounding DINO Tiny as an Open-Vocabulary object detector, which was pre-trained on the O365 [44], GoldG [16], and Cap4M [20] datasets.
For LoRA fine-tuning, we applied low-rank updates to the encoder layers with a rank of $16$. The optimization was performed using AdamW [26], with a learning rate of $1\times 10^{-3}$ and a weight decay of $1\times 10^{-2}$. We exclusively relied on Grounding DINO’s contrastive classification loss and localization loss, without introducing any additional loss functions. The final model comprised approximately $173$ million parameters, with only $19.6$ million LoRA parameters actively updated during training on ODinW-13. After fine-tuning on ODinW-13, only 75.65 MB of LoRA parameters were retained for storage.
Appendix F Limitations and Broader Impact
A limitation of the current work is that the merging strategies employed by DitHub for both class-specific experts ($A$ matrices) and the shared $B$ matrix could be further optimized. While the current solutions demonstrated effectiveness, their reliance on empirically derived coefficients highlights a potential for refinement towards strategies grounded in more robust theoretical principles. Moreover, when merging modules pertaining to the same class but originating from different tasks, incorporating explicit domain knowledge becomes crucial. We hypothesize that a mechanism explicitly accounting for domain information (e.g., distinguishing between RGB and thermal images) could not only enhance DitHub’s performance but also facilitate the development of novel expert module merging techniques. Despite these acknowledged limitations, DitHub achieves state-of-the-art performance, underscoring the efficacy of its expert module management approach.
We propose that our method, and more broadly Modular Deep Learning, can serve as a cornerstone for fostering more accessible Artificial Intelligence. The development of methods for training and managing libraries of expert modules could enable the granular selection and deployment of knowledge onto large, otherwise computationally intractable, pre-trained architectures. Specifically, DitHub facilitates the management of modules that enable large Open-Vocabulary object detectors to recognize specific and potentially rare classes. This specificity also contributes to enhanced model controllability, potentially enabling targeted unlearning capabilities for bias mitigation and thereby advancing towards fairer deep learning systems.
Appendix G Qualitative Results
@@
In Fig. D, we compare zero-shot Grounding DINO (left) with ZiRA (center) and DitHub (right). Our method excels at detecting challenging objects, including small ones that other models miss. Additionally, DitHub avoids false positives (ZiRA) and false negatives (Grounding DINO), demonstrating more reliable predictions. It also correctly identifies both cars (last row) despite occlusions and distractors. This improved stability stems from per-class specialization, which mitigates hallucinations and ensures smoother adaptation to domain shifts.